OpenStudy (anonymous):

Helena made trail mix and wants to list the ingredients in order of increasing weight. She used 1 5/8 lb of mixed nuts, 5/8 lb of chocolate covered raisins, and 1 2/8 lb of dried mixed fruit. If she decides to include 1 1/2 lb of sunflower seeds, which list shows the order of ingredients from least weight to greatest weight? A:seeds, nuts, raisins, fruit B: nuts, seeds, raisins, fruit C:raisins, fruit, nuts, seeds D:raisins, fruit, seeds, nuts

OpenStudy (anonymous):

1 5/8 = 13/8 1 2/8 = 10/8 1 1/2 = 3/2 = 12/8 so raisins - fruit - seeds - nuts

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np